two copper wires have their in ratios 2:3and lenght in ratios 3:4the ratio ohf resistance

Grade:10

1 Answers

Arun
6 years ago
Dear Nikhil
R = \rho L /A
now
 R1 / R2 = L1 A2 / L2 A1
 = 2 * 4 / 3*3 = 8 / 9
Regards
Arun (askIITians forum expert)